find the domain of x-2/x^2+6x-16
Answer by lwsshak3(11628)

find the domain of x-2/x^2+6x-16
set denominator=0 and solve for x-values which makes expression undefined
x^2+6x-16=0
(x+8)(x-2)=0
x=-8
or
x=2
domain: (-∞,-8) U (-8,2) U (2,∞)